# Discrete flavour and CP symmetries in light of JUNO and neutrino global fit

### v1

@@ 5.2 Lepton mixing matrix depending on two free parameters
There are three distinct two-parameter mixing patterns that can be derived from the group $\Delta(6n^{2})\rtimes H_{CP}$, as discussed in section 4.3. We perform a full $\chi^{2}$ analysis of these three cases for $\Delta(6n^{2})$ with $n\leq 4$. For the mixing pattern $U^{VIII}$, the best-fit values of the two real input parameters and the corresponding predictions for the lepton mixing observables are summarized in table 6, while the resulting allowed regions of these observables are presented in table 11. Taking into account permutations and inequivalent choices of $(\varphi_{1},\varphi_{2})$, we find 10 and 4 phenomenologically viable realizations for NO and IO, respectively. For NO, the viable patterns include $U^{VIII}_{i}$ with $i=4,5,6,7,8,9$. In particular, $U^{VIII}_{4,5,7,8}$ arise from $\Delta(54)$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{3},0)$, while $U^{VIII}_{6}$ and $U^{VIII}_{9}$ arise from $\Delta(96)$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$, $(\frac{\pi}{4},\frac{\pi}{4})$, and $(\frac{\pi}{4},\frac{\pi}{2})$. We note that for $\varphi_{2}=0$, the mixing matrix in Eq. (4.17) becomes real, implying CP-conserving Dirac and Majorana phases. For IO, only $U^{VIII}_{6}$ and $U^{VIII}_{9}$ remain viable, both obtained from $\Delta(96)$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},\frac{\pi}{4})$ or $(\frac{\pi}{4},\frac{\pi}{2})$. Thus the minimal $\Delta(6n^{2})$ group that can generate viable mixing pattern $U^{VIII}$ is $\Delta(54)$ with CP-conserving phases, and $\Delta(96)$ with CP-violating phases. Since the mixing pattern $U^{VIII}$ can also originate from $D_{n}\rtimes H_{CP}$ and the group order of $D_{n}$, namely $2n$, is smaller than that of $\Delta(6n^{2})$, we further scan the $D_{n}$ series up to $n\leq 8$. Beyond the viable solutions already found for $\Delta(6n^{2})$ with $n=3,4$ and $\varphi_{2}=0,\frac{\pi}{2}$ that can also be obtained from $D_{n}$, we identify an additional 6 viable cases for NO and 1 for IO, as listed in table 6. The extra viable NO patterns include $U^{VIII}_{1}$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{5},0)$ or $(\frac{\pi}{5},\frac{\pi}{2})$, $U^{VIII}_{4}$ and $U^{VIII}_{7}$ with $(\varphi_{1},\varphi_{2})=(\frac{2\pi}{5},0)$, and $U^{VIII}_{5}$ and $U^{VIII}_{8}$ with $(\varphi_{1},\varphi_{2})=(\frac{2\pi}{7},0)$. For IO, the additional viable case is $U^{VIII}_{1}$ with $(\varphi_{1},\varphi_{2})=(\frac{2\pi}{5},\frac{\pi}{2})$. For the case of $D_{n}\rtimes H_{CP}$, the smallest $D_{n}$ group capable of producing a viable $U^{VIII}$ mixing pattern is $D_{3}$, which predicts CP-conserving phases, whereas $D_{4}$ gives rise to nontrivial CPV phases.
@@
We show in figure 5 the likelihood profiles of the lepton observables for all viable $U^{VIII}$ mixing patterns with $\chi^{2}_{\text{min}}\leq 9$. Regarding $\sin^{2}\theta_{12}$, we find that $U^{VIII}_{1}$ exhibits a very narrow allowed region. This behaviour originates from the correlation between $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{13}$ in Eq. (4.23). For $\varphi_{1}=\frac{\pi}{5}$, the value of $\sin^{2}\theta_{12}$ is tightly restricted around $0.331$. Consequently, the pattern $U^{VIII}_{1}$ is disfavoured by the current JUNO 59.1-day data at the $2\sigma$ level and may be excluded by the projected six-year JUNO precision. In contrast, the predictions for $\sin^{2}\theta_{12}$ in the other viable patterns remain compatible with both the present and future JUNO sensitivities.
For $U^{VIII}_{6}$ and $U^{VIII}_{9}$, the value of $\sin^{2}\theta_{23}$ is also determined by $\sin^{2}\theta_{13}$ via Eq. (4.23), leading to very narrow likelihood profiles. We suppose in the discussion that the current NuFIT best fit values of $\sin^{2}\theta_{23}$ would not change visibly. We observe that $U^{VIII}_{4}$, $U^{VIII}_{8}$ and $U^{VIII}_{9}$ predict $\sin^{2}\theta_{23}<0.5$, which lies within the assumed future $3\sigma$ sensitivity of DUNE and T2HK for NO. In contrast, the patterns $U^{VIII}_{7}$, $U^{VIII}_{5}$ and $U^{VIII}_{6}$ which are obtained from $U^{VIII}_{4}$, $U^{VIII}_{8}$ and $U^{VIII}_{9}$ by exchanging the second and third rows—predict $\sin^{2}\theta_{23}>0.5$, and could therefore be incompatible with considered future DUNE and T2HK high precision data in the NO case. For IO, the predicted $\sin^{2}\theta_{23}$ of $U^{VIII}_{9}$ can be tested by the future DUNE and T2HK data while that of $U^{VIII}_{6}$ remains within the projected $3\sigma$ ranges.
@@
The second two-parameter mixing pattern obtained from $\Delta(6n^{2})\rtimes H_{CP}$ is denoted by $U^{IX}$. As shown in Eq. (4.29), four distinct realizations arise from different permutations of rows and columns. The $\chi^{2}$-minimization results for $U^{IX}_{1,2,3,4}$ are summarized in table 7. There are in total 16 and 11 viable cases for NO and IO respectively. The smallest $\Delta(6n^{2})$ group that can reproduce the experimental data is $S_{4}$ $(n=2)$. In this case, all four mixing patterns $U^{IX}_{1,2,3,4}$ accommodate the NuFIT results for NO with $(\varphi_{1},\varphi_{2})=(0,0)$, and the CPV phases are CP-conserving because the PMNS matrix is real.
@@
In figure 6, we show the predicted likelihood profile for case $U^{IX}$ with $\chi^{2}_{\text{min}}\leq 9$. We find that all selected viable patterns are compatible with the current JUNO 59.1-day data for both NO and IO. The cases $U^{IX}_{1}$ and $U^{IX}_{3}$ with $(\varphi_{1},\varphi_{2})=\left(\frac{\pi}{4},\frac{\pi}{2}\right)$ will be most likely excluded by the projected six-year JUNO precision for NO. From the plots of $\sin^{2}\theta_{23}$, we can conclude that cases $U^{IX}_{1}$ and $U^{IX}_{4}$ predict $\sin^{2}\theta_{23}<0.5$, while $U^{IX}_{2}$ and $U^{IX}_{3}$ predict $\sin^{2}\theta_{23}>0.5$, which means the latter cases can be tested by future DUNE and T2HK results for NO, and vice versa for IO. For $\delta_{CP}$, cases $U^{IX}_{1}$ and $U^{IX}_{2}$ may be tested by the assumed prospective future DUNE and T2HK results for both NO and IO.
@@
The numerical analysis for the pattern $U^{X}$ is summarized in table 7. Unlike $U^{IX}$, only 4 NO and 2 IO realizations are viable. As in the previous case, the smallest group capable of accommodating the data is $S_{4}$ $(n=2)$, where the realizations $U^{X}_{1,2}$ with $(\varphi_{1},\varphi_{2})=(0,0)$ imply CP conservation due to the reality of the PMNS matrix. Moving to the next viable group, $\Delta(96)$ $(n=4)$, both $U^{X}_{1}$ and $U^{X}_{2}$ become viable when $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$. Figure 7 illustrates the corresponding likelihood profiles. Owing to the sum rule in Eq. (4.35), the atmospheric angle in $U^{X}_{1}$ and $U^{X}_{2}$ is tightly correlated with $\sin^{2}\theta_{13}$, yielding sharply peaked distributions for $\sin^{2}\theta_{23}$. Anticipated sensitivities from DUNE and T2HK will be sufficient to exclude $U^{X}_{1}$ for NO and $U^{X}_{2}$ for IO. For the Dirac phase, only the configuration $U^{X}_{1}$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$ remain potentially compatible with the considered prospective future long-baseline data in the NO scenario, while the remaining options in both mass orderings are expected to be disfavored.
@@
Based on the numerical analysis above, we conclude that almost all two-parameter mixing patterns remain compatible with the current oscillation data, even when the JUNO 59.1-day measurement of $\sin^{2}\theta_{12}$ is taken into account. The discrimination among these patterns will rely on future long-baseline measurements of $\sin^{2}\theta_{23}$ and $\delta_{CP}$. For the mixing patterns originating from $A_{5}\rtimes H_{CP}$, the mixing matrix $U^{VI}_{6,1}$, featuring a fixed (21) entry of $\frac{1}{2\phi_{g}}$, achieves the best fit for NO data but exhibits CP-conserving phases. The pattern $U^{VI}_{1}$, with a fixed (11) entry of $\frac{\phi_{g}}{2}$, provides a good fit to the experimental data for both NO and IO and predicts non-trivial CP-violating Dirac and Majorana phases. For the patterns $U^{VII}$ arising from $\Sigma(168)\rtimes H_{CP}$, the fixed entry is $\frac{1}{2}$, and the minimal $\chi^{2}_{\rm min}$ among different permutations is obtained for $U^{VII}_{2,2}$. Regarding the three distinct patterns $U^{VIII}$, $U^{IX}$, and $U^{X}$, with fixed entries $\cos\varphi_{1}$, $\frac{1}{2}$ and $\frac{1}{\sqrt{2}}$, respectively, which originate from $\Delta(6n^{2})\rtimes H_{CP}$, the minimal viable groups for the case of CP-conserving phases are $\Delta(54)$, $S_{4}$ and $S_{4}$, respectively. To accommodate CP-violating $\delta_{CP}$, the group $\Delta(96)$ is required, allowing all three patterns to fit the experimental data. Furthermore, the $D_{3}$ and $D_{4}$ groups, which are the symmetry groups of equilateral triangle and square respectively, can also be employed to generate CP-conserved and CP-violating viable realizations of $U^{VIII}$, respectively. It is notable that the group order is drastically reduced with respect to $\Delta(54)$ and $\Delta(96)$. Notice that the flavor group $D_{5}$ can generate the viable mixing matrix $U^{VIII}_{1}$ with the fixed $(11)$ entry $\cos(\pi/5)$, leading to $\sin^{2}\theta_{12}$ tightly constrained around $0.331$ as dictated by Eq. (4.7). This region is consistent with the current JUNO data, but lies near the upper bound and will be subject to further scrutiny with improved JUNO precision.
@@ 6 Summary and conclusion
We have carried out a comprehensive assessment of the phenomenological viability of the discrete lepton flavour groups $A_{5}$, $\Sigma(168)$ and the series $\Delta(6n^{2})$ combined with gCP symmetry. The full symmetry groups $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$, and $\Delta(6n^{2})\rtimes H_{CP}$ are broken to non-trivial residual subgroups $\mathcal{G}_{\ell}=G_{\ell}\rtimes H^{\ell}_{CP}$ and $\mathcal{G}_{\nu}=G_{\nu}\times H^{\nu}_{CP}$ in the charged lepton and neutrino sectors, respectively. The PMNS matrix is completely determined by these residual symmetries, no assumptions about the underlying dynamical origin of symmetry breaking are required. Note that if the residual flavour symmetry $G_{\ell}$ distinguishes the three families, the inclusion of gCP symmetry $H^{\ell}_{CP}$ would not add any information as far as lepton mixing is concerned [51, 52].
@@
Assuming that the three lepton doublets transform as a faithful triplet of the flavour group, we systematically analyze all possible residual symmetries and the resulting predictions for the lepton mixing matrix. These two classes of breaking patterns lead to qualitatively different parametric structures of the PMNS matrix. Under pattern (i), the PMNS matrix depends on a single real parameter $\theta\in[0,\pi)$ and one of its columns is completely fixed. Under pattern (ii), the PMNS matrix depends on two real parameters $(\theta_{\ell},\theta_{\nu})$ in $0,\pi)$ and one matrix element is fixed. A third possibility, with $\mathcal{G}_{\ell}=Z_{2}$ and $\mathcal{G}_{\nu}=K_{4}\times H^{\nu}_{CP}$, also yields a two–parameter mixing matrix but with one row fixed. This option is incompatible with the NuFIT data and is therefore discarded; see Appendix C.
First of all, we performed a statistical test of the sum-rule predictions among the mixing angles for all patterns obtained in approaches (i) and (ii), using the latest global neutrino oscillation data from NuFIT-v6.0 [6]. Our analysis selects the mixing patterns whose predictions satisfy the following requirement: the three lepton mixing angles $(\theta_{13},\theta_{12},\theta_{23})$ and the Dirac CPV phase $\delta_{CP}$ must all lie within their experimentally allowed $3\sigma$ ranges [6]. In approach (i), all mixing patterns that remain consistent with the measured mixing angles and the $\delta_{CP}$ at the $3\sigma$ level [6] can be classified into 5 distinct cases, up to permutations of rows and columns. The corresponding mixing matrices are
@@
Notably, $U^{I}$ and $U^{II}$ arise from the breaking of $\Delta(6n^{2})$ flavour symmetries combined with CP. Furthermore, $U^{II}$ can also be realized by breaking $\Delta(3n^{2})\rtimes H_{CP}$ [78] if nontrivial singlets are absent from the model. The mixing matrices $U^{III}$ and $U^{IV}$ are minimally realized using the alternating group $A_{5}$ combined with CP symmetry, while $U^{V}$ is reproduced by the flavour symmetry group $\Sigma(168)\rtimes H_{CP}$. In the breaking approach (ii), our statistical analysis identifies 11 viable breaking patterns. Among them, six arise from the breaking of $A_{5}\rtimes H_{CP}$, yielding the PMNS matrices $U^{VI}_{i}$ with $i=1,\dots,6$ in Eq. (4.6). Two additional patterns originate from $\Sigma(168)\rtimes H_{CP}$, corresponding to $U^{VII}_{1}$ and $U^{VII}_{2}$ in Eq. (4.15). The remaining three patterns stem from $\Delta(6n^{2})\rtimes H_{CP}$ and lead to the mixing matrices $U^{VIII}$, $U^{IX}$, and $U^{X}$, shown in Eq. (4.17), Eq. (4.26), and Eq. (4.30), respectively. Furthermore, when the left-handed lepton doublets transform as a direct sum of a singlet and a doublet, the mixing matrix $U^{VIII}$ in Eq. (4.17) can also be realized through the breaking of the dihedral group $D_{n}$ combined with generalized CP symmetry. We summarize the fixed columns for the one-parameter mixing patterns $U^{I}$, $U^{II}$, $U^{III}$, $U^{IV}$ and $U^{V}$, and the fixed entries for the two-parameter mixing patterns $U^{VI}$, $U^{VII}$, $U^{VIII}$, $U^{IX}$ and $U^{X}$ in table 8.

### v2

@@ 5.2 Lepton mixing matrix depending on two free parameters
There are three distinct two-parameter mixing patterns that can be derived from the group $\Delta(6n^{2})\rtimes H_{CP}$, as discussed in section 4.3. We perform a full $\chi^{2}$ analysis of these three cases for $\Delta(6n^{2})$ with $n\leq 4$. For the mixing pattern $U^{VIII}$, the best-fit values of the two real input parameters and the corresponding predictions for the lepton mixing observables are summarized in table 6, while the resulting allowed regions of these observables are presented in table 11. Taking into account permutations and inequivalent choices of $(\varphi_{1},\varphi_{2})$, we find 10 and 4 phenomenologically viable realizations for the NO and IO mass spectra, respectively. For NO, the viable patterns include $U^{VIII}_{i}$ with $i=4,5,6,7,8,9$. In particular, $U^{VIII}_{4,5,7,8}$ arise from $\Delta(54)$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{3},0)$, while $U^{VIII}_{6}$ and $U^{VIII}_{9}$ arise from $\Delta(96)$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$, $(\frac{\pi}{4},\frac{\pi}{4})$, and $(\frac{\pi}{4},\frac{\pi}{2})$. We note that for $\varphi_{2}=0$, the mixing matrix in Eq. (4.17) becomes real, implying CP-conserving Dirac and Majorana phases. For IO, only $U^{VIII}_{6}$ and $U^{VIII}_{9}$ remain viable, both obtained from $\Delta(96)$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},\frac{\pi}{4})$ or $(\frac{\pi}{4},\frac{\pi}{2})$. Thus the minimal $\Delta(6n^{2})$ group that can generate viable mixing pattern $U^{VIII}$ is $\Delta(54)$ with CP-conserving phases, and $\Delta(96)$ with CP-violating phases. Since the mixing pattern $U^{VIII}$ can also originate from $D_{n}\rtimes H_{CP}$ and the group order of $D_{n}$, namely $2n$, is smaller than that of $\Delta(6n^{2})$, we further scan the $D_{n}$ series up to $n\leq 8$. Beyond the viable solutions already found for $\Delta(6n^{2})$ with $n=3,4$ and $\varphi_{2}=0,\frac{\pi}{2}$ that can also be obtained from $D_{n}$, we identify an additional 6 viable cases for NO and 1 for IO, as listed in table 6. The extra viable NO patterns include $U^{VIII}_{1}$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{5},0)$ or $(\frac{\pi}{5},\frac{\pi}{2})$, $U^{VIII}_{4}$ and $U^{VIII}_{7}$ with $(\varphi_{1},\varphi_{2})=(\frac{2\pi}{5},0)$, and $U^{VIII}_{5}$ and $U^{VIII}_{8}$ with $(\varphi_{1},\varphi_{2})=(\frac{2\pi}{7},0)$. For IO, the additional viable case is $U^{VIII}_{1}$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{5},\frac{\pi}{2})$. For the case of $D_{n}\rtimes H_{CP}$, the smallest $D_{n}$ group capable of producing a viable $U^{VIII}$ mixing pattern is $D_{3}$, which predicts CP-conserving phases, whereas $D_{4}$ gives rise to nontrivial CP-violating phases, with $\delta_{CP}$ having values close to $3\pi/2$.
@@
We show in figure 5 the likelihood profiles of the lepton observables for all viable $U^{VIII}$ mixing patterns with $\chi^{2}_{\text{min}}\leq 9$. Regarding $\sin^{2}\theta_{12}$, we find that the pattern $U^{VIII}_{1}$ exhibits a very narrow allowed region. This behaviour originates from the correlation between $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{13}$ in Eq. (4.23). For $\varphi_{1}=\frac{\pi}{5}$, the value of $\sin^{2}\theta_{12}$ is tightly restricted around $0.331$. Consequently, the pattern $U^{VIII}_{1}$ is disfavoured by the current JUNO 59.1-day data at the $2\sigma$ level and may be excluded by JUNO given its projected six-year precision. In contrast, the predictions for $\sin^{2}\theta_{12}$ in the other viable patterns remain compatible with both the present and future prospective JUNO precision data.
For $U^{VIII}_{6}$ and $U^{VIII}_{9}$, the value of $\sin^{2}\theta_{23}$ is also determined by $\sin^{2}\theta_{13}$ via Eq. (4.23), leading to very narrow likelihood profiles. We suppose in the discussion that the current NuFIT best fit values of $\sin^{2}\theta_{23}$ would not change significantly. We observe that $U^{VIII}_{4}$, $U^{VIII}_{8}$ and $U^{VIII}_{9}$ predict $\sin^{2}\theta_{23}<0.5$, which lies within the assumed future $3\sigma$ sensitivity of DUNE and T2HK for NO. In contrast, the patterns $U^{VIII}_{7}$, $U^{VIII}_{5}$ and $U^{VIII}_{6}$ which are obtained from $U^{VIII}_{4}$, $U^{VIII}_{8}$ and $U^{VIII}_{9}$ by exchanging the second and third rows—predict $\sin^{2}\theta_{23}>0.5$, and could therefore be incompatible with considered future DUNE and T2HK high precision data in the NO case. For IO, the predicted $\sin^{2}\theta_{23}$ of $U^{VIII}_{9}$ can be tested by the future DUNE and T2HK data while that of $U^{VIII}_{6}$ remains within the projected $3\sigma$ ranges.
@@
The second two-parameter mixing pattern obtained from $\Delta(6n^{2})\rtimes H_{CP}$ is denoted by $U^{IX}$. As shown in Eq. (4.29), four distinct realizations arise from different permutations of rows and columns. The $\chi^{2}$-minimization results for $U^{IX}_{1,2,3,4}$ are summarized in table 7. There are in total 16 and 11 viable cases for NO and IO respectively. The smallest $\Delta(6n^{2})$ group that can reproduce the experimental data is $S_{4}$ $(n=2)$. In this case, all four mixing patterns $U^{IX}_{1,2,3,4}$ accommodate the NuFIT results for NO with $(\varphi_{1},\varphi_{2})=(0,0)$, and the CPV phases are CP-conserving because the PMNS matrix is real, with $\delta_{CP}=0$ or $\pi$.
@@
In figure 6, we show the predicted likelihood profile for case $U^{IX}$ with $\chi^{2}_{\text{min}}\leq 9$. We find that all selected viable patterns are compatible with the current JUNO 59.1-day data for both NO and IO. The cases $U^{IX}_{1}$ and $U^{IX}_{3}$ with $(\varphi_{1},\varphi_{2})=\left(\frac{\pi}{4},\frac{\pi}{2}\right)$ will be most likely excluded for NO by the projected six-year JUNO precision data. From the plots of $\sin^{2}\theta_{23}$, we can conclude that the patterns $U^{IX}_{1}$ and $U^{IX}_{4}$ predict $\sin^{2}\theta_{23}<0.5$, while $U^{IX}_{2}$ and $U^{IX}_{3}$ predict $\sin^{2}\theta_{23}>0.5$. Both sets of patterns can be tested for both the NO and IO scenarios by the future DUNE and T2HK prospective results on $\sin^{2}\theta_{23}$ and $\delta_{CP}$.
@@
The numerical analysis for the pattern $U^{X}$ is summarized in table 7. Unlike $U^{IX}$, only 4 NO and 2 IO realizations are viable. As in the previous case, the smallest group capable of accommodating the data is $S_{4}\cong\Delta(6\cdot 2^{2})$, where the realizations $U^{X}_{1,2}$ with $(\varphi_{1},\varphi_{2})=(0,0)$ imply CP conservation due to the reality of the PMNS matrix, with $\delta_{CP}=\pi$, which for NO is not excluded by the current data. Moving to the next viable group, $\Delta(96)=\Delta(6\cdot 4^{2})$, the predictions of both the $U^{X}_{1}$ and $U^{X}_{2}$ patterns are compatible with the current data for $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$. Figure 7 illustrates the corresponding likelihood profiles. Owing to the sum rule in Eq. (4.35), the atmospheric angle in $U^{X}_{1}$ and $U^{X}_{2}$ is tightly correlated with $\sin^{2}\theta_{13}$, yielding sharply peaked distributions for $\sin^{2}\theta_{23}$. The prospective sensitivities T2HK and DUNE would be sufficient to possibly exclude the pattern $U^{X}_{1}$ for NO and the pattern $U^{X}_{2}$ for IO. For the Dirac phase, only the configuration $U^{X}_{1}$ with $(\varphi_{1},\varphi_{2})=(\frac{\pi}{4},0)$ remain potentially compatible with the considered prospective future long baseline data in the NO scenario, while the remaining options in both mass orderings are expected to be disfavored.
@@
Based on the numerical analysis above, we conclude that almost all two-parameter mixing patterns remain compatible with the current oscillation data, even when the JUNO 59.1-day measurement of $\sin^{2}\theta_{12}$ is taken into account. The discrimination between these patterns will rely on future long baseline measurements of $\sin^{2}\theta_{23}$ and $\delta_{CP}$. For the mixing patterns originating from $A_{5}\rtimes H_{CP}$, the mixing matrix $U^{VI}_{6,1}$, featuring a fixed (21) entry of $\frac{1}{2\phi_{g}}$, gives the best fit for NO data, while predicting CP-conserving phases. The pattern $U^{VI}_{1}$, with a fixed (11) entry of $\frac{\phi_{g}}{2}$, provides a good fit to the experimental data for both NO and IO spectra and predicts non-trivial CP-violating Dirac and Majorana phases. For the patterns $U^{VII}$ arising from $\Sigma(168)\rtimes H_{CP}$, the fixed entry is $\frac{1}{2}$, and the minimal $\chi^{2}_{\rm min}$ among different permutations is obtained for $U^{VII}_{2,2}$. Regarding the three distinct patterns $U^{VIII}$, $U^{IX}$, and $U^{X}$, with fixed entries $\cos\varphi_{1}$, $\frac{1}{2}$ and $\frac{1}{\sqrt{2}}$, respectively, which originate from $\Delta(6n^{2})\rtimes H_{CP}$, the minimal viable groups for the case of CP-conserving phases are $\Delta(54)$, $S_{4}$ and $S_{4}$, respectively. To accommodate CP-violating $\delta_{CP}$, the group $\Delta(96)$ is required, allowing all three patterns to fit the experimental data. Furthermore, the $D_{3}$ and $D_{4}$ groups, which are the symmetry groups respectively of the equilateral triangle and the square, can also be employed to generate CP-conserved and CP-violating viable realizations of $U^{VIII}$, respectively. It is notable that the group order is drastically reduced with respect to $\Delta(54)$ and $\Delta(96)$. We notice that the flavor group $D_{5}$ can generate the viable mixing matrix $U^{VIII}_{1}$ with the fixed $(11)$ entry $\cos(\pi/5)$, leading to $\sin^{2}\theta_{12}$ tightly constrained around $0.331$ as dictated by Eq. (4.7). This region is consistent with the current JUNO data, but lies near the upper bound and will be subject to further scrutiny with improved JUNO precision.
@@ 6 Summary and conclusion
We have carried out a comprehensive assessment of the phenomenological viability of the lepton flavour models based on the non-Abelian discrete groups $A_{5}$, $\Sigma(168)$ and the series $\Delta(6n^{2})$ combined with gCP symmetry. The full symmetry groups $A_{5}\rtimes H_{CP}$, $\Sigma(168)\rtimes H_{CP}$, and $\Delta(6n^{2})\rtimes H_{CP}$ are broken to non-trivial residual subgroups $\mathcal{G}_{\ell}=G_{\ell}\rtimes H^{\ell}_{CP}$ and $\mathcal{G}_{\nu}=G_{\nu}\times H^{\nu}_{CP}$ in the charged lepton and neutrino sectors, respectively. The PMNS matrix is completely determined by these residual symmetries, no assumptions about the underlying dynamical origin of symmetry breaking are required. We note that if the residual flavour symmetry $G_{\ell}$ distinguishes the three families, the inclusion of the gCP symmetry $H^{\ell}_{CP}$ would not add any information on the lepton mixing [54, 55]. It is remarkable that we can reach sharp predictions for both Dirac CP violation phase $\delta_{CP}$ and Majorana CP phases $\alpha_{21}$, $\alpha_{31}$ due to the presence of gCP symmetry.
@@
Assuming that the three lepton doublets transform as a faithful triplet of the flavour group, we systematically analyze all possible residual symmetries and the resulting predictions for the lepton mixing matrix. The indicated two classes of breaking patterns lead to qualitatively different parametric structures of the PMNS matrix. Under pattern (i), the PMNS matrix depends on a single real parameter $\theta\in[0,\pi)$ and one of its columns is completely fixed. Under pattern (ii), the PMNS matrix depends on two real parameters $(\theta_{\ell},\theta_{\nu})$ in $0,\pi)$ and one matrix element is fixed. A third possibility, with $\mathcal{G}_{\ell}=Z_{2}$ and $\mathcal{G}_{\nu}=K_{4}\times H^{\nu}_{CP}$, also yields a two–parameter mixing matrix but with one row fixed. This option is incompatible with the NuFIT data and is therefore discarded; see Appendix C.
First of all, we performed a statistical test of the predicted correlations, in the form of sum rules, between the mixing angles for all patterns obtained in approaches (i) and (ii), using the latest global neutrino oscillation data from NuFIT-v6.0 [9]. Our analysis selects the mixing patterns whose predictions satisfy the following requirement: the three lepton mixing angles $(\theta_{13},\theta_{12},\theta_{23})$ and the Dirac CPV phase $\delta_{CP}$ must all lie within their experimentally allowed $3\sigma$ ranges [9]. In approach (i), all mixing patterns that remain consistent with the measured mixing angles and the $\delta_{CP}$ at the $3\sigma$ level [9] can be classified into 5 distinct cases, up to permutations of rows and columns. The corresponding mixing matrices are
@@
Notably, $U^{I}$ and $U^{II}$ arise from the breaking of $\Delta(6n^{2})$ flavour symmetries combined with the generalized CP (gCP) symmetry. Furthermore, $U^{II}$ can also be realized by breaking $\Delta(3n^{2})\rtimes H_{CP}$ [81] if nontrivial singlets are absent from the model. The mixing matrices $U^{III}$ and $U^{IV}$ are minimally realized using the alternating group $A_{5}$ in combination with the gCP symmetry, while $U^{V}$ is reproduced by the flavour symmetry group $\Sigma(168)\rtimes H_{CP}$. In the breaking approach (ii), our statistical analysis identifies 11 viable breaking patterns. Among them, six arise from the breaking of $A_{5}\rtimes H_{CP}$, yielding the PMNS matrices $U^{VI}_{i}$ with $i=1,\dots,6$ in Eq. (4.6). Two additional patterns originate from $\Sigma(168)\rtimes H_{CP}$, corresponding to $U^{VII}_{1}$ and $U^{VII}_{2}$ in Eq. (4.15). The remaining three patterns stem from $\Delta(6n^{2})\rtimes H_{CP}$ and lead to the mixing matrices $U^{VIII}$, $U^{IX}$, and $U^{X}$, shown in Eq. (4.17), Eq. (4.26), and Eq. (4.30), respectively. Furthermore, when the left-handed lepton doublets transform as a direct sum of a singlet and a doublet, the mixing matrix $U^{VIII}$ in Eq. (4.17) can also be realized through the breaking of the dihedral group $D_{n}$ combined with the gCP symmetry. We list the fixed columns for the one-parameter mixing patterns $U^{I}$, $U^{II}$, $U^{III}$, $U^{IV}$ and $U^{V}$, and the fixed entries for the two-parameter mixing patterns $U^{VI}$, $U^{VII}$, $U^{VIII}$, $U^{IX}$ and $U^{X}$ in table 8.